Tuesday 10/30/12

9:30 A – 7:30P, 8:30P – 11:50 P


  • Flip collection plates.
  • # Save MTDs to plate into new bottles.  All in one, need high density.
  • Collect virgin Espl/sim, cross to TM2/TM6 males.
  • Collect virgin Act5-gal4
  • Make new AJ plates
  • Fix embryos: regular 25 min fix in 8% FA: MTDxPsc 1, MTDxPc 3, MTDxScm, and MTD straight up.
  • PM virgin collection, Act5-Gal4,
  • PM virgin collection Espl/sim, cross to TM2/TM6
  • New collection cages:
    •  MTDxScm (small cage #2)
    • sim[D] large cage
  • Fix embryos PM. regular 25 min fix in 8% FA MTDxPsc 1, MTDxPc 3, MTDxScm,
  • Short 15 min fix in 8% FA: MTD straight up.
    • Moved to EtOH then PBT washes.  Blocking prior to O/N incubation in anti-Pc 1:100.
    • 5 min post-fix tomorrow then move back to MeOH for 1 week plus.  Previous sample was stored in MeOH at -20C for ~20hrs, but RNA probes didn’t label at all.  Meanwhile identical probes set used on yw labeled in parallel worked great.

In situs cell culture  day 2

  • Rigerous 55C washout of probe (hopefully don’t lose all the attached cells).
  • PBS rinse, Antibody labeling,
  • [spreadsheet 0AjSqkxgziU1YdG90a29UaU5ua1FLUXZ5LTFmM1pMSlE 602 202 sheet=9]
  • Primary antibodies started, 4P
  • Check staining on Turnkey: still substantial background in RNA probe channels.  No probe control is plenty clean, so it is definetely a probe sticking thing.
  • Potential nascent foci in several cells.  hard to tell.  Background might be substantially lower than previous RNA stains, or it might be a sensitivity of the microscope (will validate on confocal).   Maybe try doing all hot washes at 65?  No beacon bright nascent transcripts though.  Antibodies get in fine though so RNA probes should also make it?
  • check on confocal at 8P tomorrow.
DNA FISH probes
  • stop O/N reaction
  • photograph test gel
  • isolate DNA probes,  (now at -80C, precipitating).  
  • spin down, rota-evap, and resuspend in DFISH hybe
  • Revisions back from ML. 
  • Review these. Discussion of 4x snail got reverted back to old version.  Fixed this, wrote to Mike aobut substantive change (was supposed to just make changes and submit, but should probably get the okay since this is a whole paragraph different rather than minor wording).  
  • Current versions CR response r5, snail kinetics r4.
  • Revisions back from ML (snail kinetics r5).  Last set of minor tweaks reviewed.  Fix all gene names (italics) and protein names (capital non-italics).   Final version is snail_kinetics r6.
  • Go ahead for Submit to Editors Office (try tomorrow, it’s too late now I might screw it up).
  • Rob to investigate details of ICAM proposal and to take first shot at moving NSF proposal into word doc format.  
  • working on enhancer section of BJ review.
Cell Culture
  • Joe has adherent BG3s.  BG3 protocol 
  • get shaker from Alec.  Setup in equipment room.  
  • Passage cells into 2 new mid-size flasks (looks good). 
  • from small flask plate cells into new 8 well chamber (didn’t detach cells well enough, density is moderately low).  
  • # Get Triptan blue, check viability of sample of cells. (powder in glass cabinet downstairs)


  • moving Fall 2011 confocal data to Alistair5
  • # delete copy on data drive D.
  • Started new staining record for cultured cells
  • schedule meeting for next week with Brian and Fred


This entry was posted in Summaries and tagged , , . Bookmark the permalink.